OpenStudy (lataesiaaa):

A country's population in 1993 was 94 million. In 1999 it was 99 million. Estimate the population in 2005 using the exponential growth formula. Round your answer to the nearest million. P = Aekt

OpenStudy (freckles):

Let 1993 correspond to t=0 so 1994 corresponds to t=1 and 1995 corresponds to t=2 and 1996 corresponds to t=3 what does 1999 correspond to in terms of t? and what does 2005 correspond to in terms of t ?

OpenStudy (lataesiaaa):

1999 corresponds to t=5 2005 corresponds to t=11 ??

OpenStudy (freckles):

hmmm... I think it should be t=6 and t=12 since 1999-1993=6 and 2005-1993=12 anyways... we are going to let P be in millions so we have the following two points given for (t,P) (0,94) and (6,99) and we are asked to find P when we have t=12

OpenStudy (freckles):

so we have the equation P=A e^(kt) we can find A pretty easily using the first point (0,94)

OpenStudy (freckles):

(t,P)=(0,94) put in 0 for t and 94 for P and solve for A.

OpenStudy (lataesiaaa):

okk holdon

OpenStudy (lataesiaaa):

lol nvm i just dont get this.

OpenStudy (freckles):

\[P=Ae^{kt} \\ (t,P)=(0,94) \\ 94=Ae^{k \cdot 0}\] k*0 is...

OpenStudy (lataesiaaa):

is that the answer?

OpenStudy (freckles):

no

OpenStudy (lataesiaaa):

okay what do i have to do ?

OpenStudy (freckles):

We need to find A and k first before we can find what P corresponds to t=12

OpenStudy (freckles):

we are using the point (0,94) to find A and then we are going to use (6,99) to find k

OpenStudy (freckles):

\[P=Ae^{kt} \\ (t,P)=(0,94) \\ 94=Ae^{k \cdot 0} \] can you find A here?

OpenStudy (freckles):

e^(k*0) is a very pretty number

OpenStudy (freckles):

what number is it

OpenStudy (freckles):

assuming a isn't 0 then a^0=1 so e^(k*0)=e^0=?

OpenStudy (lataesiaaa):

lmfao thanks for your help but this is so hard and konfusing, so im just not gone take the test ..... :/

OpenStudy (freckles):

any number except 0 to the 0 power is just 1

OpenStudy (freckles):

\[P=Ae^{kt} \\ (t,P)=(0,94) \\ 94=Ae^{k \cdot 0} \\ 94=A(1) \\ 94=A \] if you decide not to give up try to find k in \[P=94 e^{kt}\] using the other point which was (6,99) replace t with 6 and P with 99 and solve for k

OpenStudy (freckles):

after that you will be already to figure out what happens in 2005